# Causal Classification of Pathological Misner-Type Spacetimes

### v4

@@ Acknowledgement.
Parts of this work were completed during my time as a member of Richard Schoen’s research group at the University of California, Irvine, and at the Simons Center for Geometry and Physics, Stony Brook University. However, the research presented here originated from a project conducted under the supervision of Kip S. Thorne during 2016-2018. I am profoundly grateful to Kip S. Thorne for his invaluable advice, guidance, supportive remarks, and encouragement throughout the course of this work. I would also like to express my deep appreciation for his inspiring spirit of intellectual adventure in research and the insightful discussions, particularly on Misner space, which have been truly indispensable to this work. Furthermore, I am very grateful to the referees for their thorough and constructive feedback.

### v5

@@ Acknowledgement.
Parts of this work were completed during my time as a member of Richard Schoen’s research group at the University of California, Irvine, and at the Simons Center for Geometry and Physics, Stony Brook University. The research presented here originated from a project conducted under the supervision of Kip S. Thorne during 2016–2018. I am profoundly grateful to Kip S. Thorne for his invaluable advice, guidance, supportive remarks, and encouragement throughout the course of this work. I would also like to express my deep appreciation for his inspiring spirit of intellectual adventure in research and the insightful discussions, particularly on Misner space, which have been truly indispensable to this work. I also wish to thank the referees for their careful reading and constructive comments, which have helped improve the manuscript.